I need help in order to understand this problem. If anyone can help I would be very grateful. Thank you very much for your assistance.

1000 raffle tickets are sold at $3 each. One grand prize of $100 and two consolation prizes of $75 will be awarded. Find Jake’s expectation if he purchases one ticket.
A) -$2.75 B) -$3.00 C) $0.24 D) $2.75 E) $3.00


Answer by stanbon(75887)

1000 raffle tickets are sold at $3 each. One grand prize of $100 and two consolation prizes of $75 will be awarded. Find Jake’s expectation if he purchases one ticket.
A) -$2.75 B) -$3.00 C) $0.24 D) $2.75 E) $3.00
----------------
Let the random variable represent Jake's earnings.
Values of the random variable are 97, 72, -3
Corresponding probabilities are (1/1000), (2/1000), (997/1000)
--------------------------
Jake's expected earnings are 97(1/100)+72(2/1000) - 3(997/1000
= (97 + 144 - 2991)/1000 = -2750/1000 = -$2.75
======================================================
Cheers,
Stan H.